OpenStudy (anonymous):

What are some Ways to make 1000. out of 11 and 7?

OpenStudy (anonymous):

82 x 11 + 11 x 54

OpenStudy (anonymous):

do you mean with the sum of multiples of 11 and multiples of 7?

OpenStudy (anonymous):

yes

OpenStudy (mani_jha):

You mean, by multiplication or addition? Well, I have got an idea. 11x+7y=1000 This is a straight line equation. Plot it in a graph. See if any integer coordinates lie on the line. If you find such a point, that is a solution for (x.y) Do you get my idea?

OpenStudy (anonymous):

i also think he means integer x and y's because there was a similar problem earlier...

OpenStudy (anonymous):

No guys I think he means something like( 89 x 11)+(3 x 7)

OpenStudy (anonymous):

What he said^

OpenStudy (mani_jha):

Yes, (89,3) is a solution to the equation I posted above.

OpenStudy (mani_jha):

Just plot it. By inspection, you could find it. Keep putting y=1,y=2 etc and see if an integral value of x results. You will get so if you put y=3

OpenStudy (anonymous):

im in 8th grade. can you please tell me something that makes sense??

OpenStudy (anonymous):

so we are talking about an equation that looks like x7 + y11 = 1000 aren't we?

OpenStudy (anonymous):

yes. but i have no idea what the variables could be

OpenStudy (anonymous):

If you start with x = 0 you are left with y11 = 1000 divide by 11 on each side y = 1000/11

OpenStudy (anonymous):

Then just any number that you pick for x is your x subtract your x*7 from 1000 and then solve for your next y

OpenStudy (anonymous):

so x=2 gives you (2)(7) +y11 = 1000 = 14 + y11 = 1000 = y11 = 986 y=986/11

OpenStudy (anonymous):

yo can you use 8th grade vocabulary